# BARI-TH/25-772

### v1

@@ 4.1.1 Exclusive lepton flavour conserving $b\to s\ell^{+}\ell^{-}$ modes
Table 1: Ranges of branching fractions of LFV modes in the scenario A of the ABCD model. The first three rows correspond to $M_{Z^{\prime}}=1$ TeV, the last three rows to $M_{Z^{\prime}}=3$ TeV. ${\cal B}^{(1)}$ is obtained after imposing that the corresponding lepton flavour conserving mode lies in the experimental range; ${\cal B}^{(2)}$ is obtained after imposing also the constraint from the upper limit of ${\cal B}(\mu^{-}\to e^{-}\gamma)$. The experimental upper bounds (at $90\%$ C.L.) are quoted in the last column.
@@
Figure 14: Correlation between the branching fractions ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded vertical band corresponds to the experimental world average [24].
@@
4.1.2 Exclusive lepton flavour violating modes $b\to s\ell_{1}^{-}\ell_{2}^{+}$ (with $\ell_{1}\neq\ell_{2}$)
A feature of the ABCD model is that correlations are generated between lepton flavour conserving and lepton flavour violating modes. In Table 1 we quote the predictions for a set of LFV modes, for $M_{Z^{\prime}}=1$ TeV and $3$ TeV. The predictions can be further constrained requiring agreement of the rates of corresponding LFC modes with experiment. In the Table we quote the unconstrained result and the one constrained by the corresponding LFC process. Visible rates for the LFV modes are possible, even though they are much smaller than the present experimental upper bounds. [^9] Recent experimental investigations of $\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-}$ are described in [76, 77]. However, the upper limits are expected to be sensibly improved in the near future.
Figure 14 shows the correlation between the branching ratios of the LFC $B_{s}\to\mu^{+}\mu^{-}$ and LFV $B_{s}\to\tau^{+}\mu^{-}$ modes for the two values of $M_{Z^{\prime}}$. The experimental upper bound for ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ is quoted in Table 1. In the ABCD model the branching ratio of the LFV $\tau^{+}\mu^{-}$ mode can be comparable to that of the LFC $\mu^{+}\mu^{-}$ mode. Indeed, the decay to same flavour leptons is determined only by the $C_{10}$ term, while for $\ell_{1}\neq\ell_{2}$ also $C_{9}$ contributes. Moreover, the helicity suppression in the $\ell_{1}=\ell_{2}=\mu$ mode is lifted in the $\tau^{+}\mu^{-}$ mode producing a sizable enhancement.
While for the LFC modes the results do not sensibly change varying $M_{Z^{\prime}}$ from 1 TeV to 3 TeV, this is not true for the LFV modes. The reason is that ABCD does not predict large deviations from SM: increasing the $Z^{\prime}$ mass, the SM results are approached. For modes allowed in SM, the deviations are small even for small values of $M_{Z^{\prime}}$. On the other hand, for modes forbidden in SM, increasing $M_{Z^{\prime}}$ results in a sensible decrease of the branching fractions. The same observation holds for ${\bar{B}}^{0}\to{\bar{K}}^{*0}\tau^{+}\mu^{-}$: the branching fraction predicted in the ABCD model is quoted in Table 1 both without and with the constraint given by the measured ${\cal B}({\bar{B}}^{0}\to{\bar{K}}^{*0}\mu^{+}\mu^{-})$. A further constraint given by the upper bound of ${\cal B}(\mu^{-}\to e^{-}\gamma)$ further reduces the range for this rate. Predictions for $B_{s}\to\phi\tau^{+}\mu^{-}$ are presented in the same Table 1 and in Fig.16.
Figure 16: Correlation between ¯${\bar{\cal B}}(B_{s}\to\phi\mu^{+}\mu^{-})$ and ¯${\bar{\cal B}}(B_{s}\to\phi\mu^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). Same color code as in Fig. 13.
@@ 4.2 Correlations among lepton flavour violating processes in the quark and lepton sectors
Figure 18: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}\gamma)$ for $M_{Z^{\prime}}=1$ TeV (top panels) and $M_{Z^{\prime}}=3$ TeV (bottom panels). The right plots are a zoom of the left ones on the region selected imposing the experimental upper bound for ${\cal B}(\mu^{-}\to e^{-}\gamma)$ [78]. The light yellow points are selected requiring that the LFC ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ agrees within $1\sigma$ with experiment.
The $Z^{\prime}$ couplings to leptons and to quarks involve the same couple of parameters $\epsilon_{1}$ and $\epsilon_{2}$. The consequence is that correlations can be established among observables in the different sectors. Important cases are the correlations among the LFV $b\to s\ell_{1}^{+}\ell_{2}^{-}$ modes and the LFV leptonic decays $\mu^{-}\to e^{-}\gamma$ and $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$.
[…]
The correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}\gamma)$ is displayed in Fig. 18, and similar remarks hold. In particular, the bound on ${\cal B}(\mu^{-}\to e^{-}\gamma)$ [78] reduces the allowed range for ${\cal B}(B\to K^{*}\tau^{+}\mu^{-})$, see Table 1. The same occurs for $B_{s}\to\phi\tau^{+}\mu^{-}$, as we report in the same Table.
As for the correlations involving ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})$, the experimental upper bound ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})<2.1\times 10^{-8}$ [24] must be taken into account, however it does not provide significant constraints for $B_{s}$ and $B$ modes, as it can be inferred from Figs. 19. Accounting for the measurement of the rate of $B_{s}\to\mu^{+}\mu^{-}$, the largest value for ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})$ is about $1.2\times 10^{-8}$, a relevant result when planning new facilities.
[…]
@@
Figure 20: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and $\Delta a_{\mu}$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The light yellow points are selected requiring that the LFC ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ agrees at $1\sigma$ with experiment.
@@
A final remark concerns the $Z^{\prime}$ contribution to the muon anomalous magnetic moment and to its correlation with other observables. The results for $\Delta a_{\mu}$ vs ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ are displayed in Fig. 20. $\Delta a_{\mu}$ turns out to be negative, and it reaches at most $\Delta a_{\mu}\sim-10^{-12}$, a size similar to subleading hadron contributions, such as the hadronic light-by-light ones [11].

### v2

@@ 4.1.1 Exclusive lepton flavour conserving $b\to s\ell^{+}\ell^{-}$ modes
Table 1: Ranges of branching fractions of LFV $B$, $B_{s}$ decays in the scenario A of the ABCD model. The first three rows correspond to $M_{Z^{\prime}}=1$ TeV, the last three rows to $M_{Z^{\prime}}=3$ TeV. The four quoted ranges are obtained after imposing the constraints from the upper bounds on LFV charged lepton decay modes specified in Table 2: ${\cal B}^{(1)}$ from $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, ${\cal B}^{(2)}$ from ${\cal B}(\mu^{-}\to e^{-}\gamma)$, ${\cal B}^{(3)}$ from $\mu^{-}\to e^{-}e^{+}e^{-}$, ${\cal B}^{(4)}$ from $\mu^{-}\to e^{-}$ conversion in titanium. The experimental upper bounds (at $90\%$ C.L.) are quoted in the last column.
@@
Figure 14: Correlation between the branching fractions ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ and ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded vertical band corresponds to the experimental world average [Navas:PDG].
@@
4.1.2 Exclusive lepton flavour violating modes $b\to s\ell_{1}^{-}\ell_{2}^{+}$ ($\ell_{1}\neq\ell_{2}$)
A feature of the ABCD model is that correlations are generated between lepton flavour conserving and lepton flavour violating modes. In Table 1 we quote the predictions for a set of LFV modes, for $M_{Z^{\prime}}=1$ TeV and $3$ TeV. The predictions can be further constrained requiring agreement of the rates of corresponding LFC modes with experiment. In the Table we quote the unconstrained result and the one constrained by the corresponding LFC process. The current experimental upper bounds for the LFV modes are also included in the Table, they are expected to be sensibly improved in the near future. [^8] Recent experimental investigations of $\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-}$ are described in [LHCb:2022wrs, Belle-II:2025hpl].
Figure 14 shows the correlation between the branching ratios of the LFC $B_{s}\to\mu^{+}\mu^{-}$ and LFV $B_{s}\to\tau^{+}\mu^{-}$ modes for the two values of $M_{Z^{\prime}}$. The experimental upper bound for ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ is quoted in Table 1. In the ABCD model the branching ratio of the LFV $\tau^{+}\mu^{-}$ mode can be comparable to that of the LFC $\mu^{+}\mu^{-}$ mode. Indeed, the decay to same flavour leptons is determined only by the $C_{10}$ term, while for $\ell_{1}\neq\ell_{2}$ also $C_{9}$ contributes. Moreover, the helicity suppression in the $\ell_{1}=\ell_{2}=\mu$ mode is lifted in the $\tau^{+}\mu^{-}$ mode producing a sizable enhancement.
While for the LFC modes the results do not sensibly change varying $M_{Z^{\prime}}$ from 1 TeV to 3 TeV, this is not true for the LFV modes, since ABCD does not predict large deviations from SM: increasing the $Z^{\prime}$ mass, the SM results are approached. For modes allowed in SM, the deviations are small even for small values of $M_{Z^{\prime}}$. On the other hand, for modes forbidden in SM, increasing $M_{Z^{\prime}}$ results in a sensible decrease of the branching fractions. The same observation holds for ${\bar{B}}^{0}\to{\bar{K}}^{*0}\tau^{+}\mu^{-}$: the branching fraction predicted in the ABCD model is quoted in Table 1 both without and with the constraint given by the measured ${\cal B}({\bar{B}}^{0}\to{\bar{K}}^{*0}\mu^{+}\mu^{-})$. As we discuss below, constraints provided by LFV leptonic decays further reduce the range for this rate. Predictions for $B_{s}\to\phi\tau^{+}\mu^{-}$ are presented in Table 1 and in Fig.16.
Figure 16: Correlation between ${\bar{\cal B}}(B_{s}\to\phi\mu^{+}\mu^{-})$ and ${\bar{\cal B}}(B_{s}\to\phi\mu^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). Same color code as in Fig. 13.
@@ 4.2 Correlations among lepton flavour violating processes in the quark and lepton sectors
Since the $Z^{\prime}$ couplings to leptons and quarks involve the same pair of parameters, $\epsilon_{1}$ and $\epsilon_{2}$, correlations can be established among observables in the different sectors. Particularly relevant are the correlations between the LFV $b\to s\ell_{1}^{+}\ell_{2}^{-}$ modes and the LFV leptonic decays $\mu^{-}\to e^{-}\gamma$, $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, $\mu^{-}\to e^{-}e^{+}e^{-}$, and the $\mu^{-}\to e^{-}$ conversion in nuclei. As we shall see, in the scenario A of the ABCD model the upper bounds on the last four channels play a hierarchical role in constraining the branching ratios of LFV $B$ and $B_{s}$ decays. Their combined effect prevents the branching fractions of LFV $B$ and $B_{s}$ decays from exceeding ${\cal O}(10^{-12})$.
It is useful to analyse the individual constraints from the four leptonic processes, which in scenarios different from scenario A depend on distinct parameters. This provides us with a guidance on possible modifications of the scenario. For instance, since $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ provides the weakest constraint, one may consider assigning a different role to the third lepton family compared to the other two: we defer such an investigation to future work. In the following we analyze the impact of $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ and $\mu^{-}\to e^{-}\gamma$, then we consider the most stringent channels $\mu^{-}\to e^{-}e^{+}e^{-}$ and $\mu^{-}\to e^{-}$ conversion in nuclei.
4.2.1 Rare $B_{(s)}$ decays versus $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ and $\mu^{-}\to e^{-}\gamma$.
@@
$\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ $2.1\times 10^{-8}$ [Navas:PDG]
$\mu^{-}\to e^{-}\gamma$ $1.5\times 10^{-13}$ [MEGnew]
$\mu^{-}\to e^{-}e^{+}e^{-}$        $1.0\times 10^{-12}$[SINDRUM:1987nra, Navas:PDG]
$\mu^{-}\to e^{-}$ conversion in $^{48}_{22}{\rm{Ti}}$     $4.3\times 10^{-12}$ [Wintz:1998rp, Kaulard:1998rb]
Table 2: Experimental upper bounds (at $90\,\%$ C.L.) for LFV charged lepton modes. The branching fraction for the $\mu^{-}\to e^{-}$ conversion is defined in Eq. (70).
@@
Figure 17: Correlation between ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}\gamma)$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The vertical line indicates the experimental upper bound ${\cal B}(\mu^{-}\to e^{-}\gamma)<1.5\times 10^{-13}\,\,(\rm{at}\,90\%\,{\rm CL})$ [MEGnew]. The light yellow points are selected requiring that the LFC ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ agrees within $1\sigma$ with experiment.
@@
Figure 18: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}\gamma)$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The vertical line corresponds to the experimental upper bound for ${\cal B}(\mu^{-}\to e^{-}\gamma)$ [MEGnew]. The light yellow points are selected requiring that the LFC ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ agrees within $1\sigma$ with experiment.
The correlations involving $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, with the experimental upper bound ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})<2.1\times 10^{-8}$ [Navas:PDG] do not provide significant constraints for $B_{s}$ and $B$ modes. Accounting for the measurement of the width of $B_{s}\to\mu^{+}\mu^{-}$, the largest value for ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})$ is about $1.2\times 10^{-8}$.
$\mu^{-}\to e^{-}\gamma$ puts more stringent limits. Figure 17 shows the correlation between ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}\gamma)$ for $M_{Z^{\prime}}=1$ TeV and $3$ TeV. The dark points show the results without imposing constraints from LFC mode, the light points the results obtained imposing that the measured ${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ is recovered. The upper bound is displayed ${\cal B}(\mu^{-}\to e^{-}\gamma)<1.5\times 10^{-13}\,\,(\rm{at}\,90\%\,{\rm CL})$ recently established by the MEG II Collaboration[MEGnew] which has improved the previous limit ${\cal B}(\mu^{-}\to e^{-}\gamma)<4.2\times 10^{-13}\,\,(\rm{at}\,90\%\,{\rm CL})$ [MEG:2016leq]. The bound reduces the room for ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$. This provides us with a clear example of the most peculiar feature of the ABCD model, the interplay between quark and lepton observables. Within the model, quite large values of ${\cal B}(\mu^{-}\to e^{-}\gamma)$ could be obtained, however, the requirement that the rate of $B_{s}\to\mu^{+}\mu^{-}$ agrees with experiment suppresses ${\cal B}(\mu^{-}\to e^{-}\gamma)$. The correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}\gamma)$ is displayed in Fig. 18, and similar remarks hold: the bound on ${\cal B}(\mu^{-}\to e^{-}\gamma)$ [MEGnew] reduces the allowed range for ${\cal B}(B\to K^{*}\tau^{+}\mu^{-})$, see Table 1. The same occurs for $B_{s}\to\phi\tau^{+}\mu^{-}$.
4.2.2 Rare $B_{(s)}$ decays versus $\mu^{-}\to e^{-}e^{+}e^{-}$ and $\mu^{-}\to e^{-}$ conversion in nuclei
The processes $\mu^{-}\to e^{-}e^{+}e^{-}$ and $\mu^{-}\to e^{-}$ conversion in nuclei play the major role in constraining LFV rare $B_{(s)}$ decays. Fig. 19 shows that the upper limit ${\cal B}(\mu^{-}\to e^{-}e^{+}e^{-})<1\times 10^{-12}\,\,(\rm{at}\,90\%\,{\rm CL})$ [SINDRUM1988, Navas:PDG] significantly reduces ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ both for $M_{Z^{\prime}}=1$ TeV and for $M_{Z^{\prime}}=3$ TeV. After imposing this bound their largest allowed values are ${\cal O}(10^{-11})$, see Table 2. Muon conversion in nuclei provides the most stringent constraint, as displayed in Fig. 20. The largest allowed value for ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ are reduced to ${\cal O}(10^{-13})$, see Table 2.
A summary of the impact of the lepton decays on ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ is in Fig. 22, where we plot their correlation when subsequent constraints from LFV charged lepton decay modes are imposed. It is instructive to examine the mutual impact of different LFV lepton decays. As an example, Fig. 21 shows the constraints imposed by the experimental upper bound for ${\cal B}(\mu^{-}\to e^{-}e^{+}e^{-})$ on ${\cal B}(\mu^{-}\to e^{-})$ and ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})$: in both cases the allowed ranges are substantially reduced.
The scrutiny of $\mu^{-}\to e^{-}e^{+}e^{-}$ and $\mu^{-}\to e^{-}$ conversion will be the focus of future experimental programs, in particular at the COMET [COMET:2018auw], Mu2e [Diociaiuti:2024stz] and Mu3e [Amarinei:2025ntv] facilities that aim at an unprecedented sensitivity. The experimental bounds used in our study are based on measurements performed more than two decades ago, and have been obtained relying on theoretical formulae and approximations that need to be reconsidered. For example, for $\mu^{-}\to e^{-}e^{+}e^{-}$ the quoted upper bound has been obtained assuming a constant decay amplitude over the entire phase space. Such assumptions will be refined in the future studies.
Figure 19: Correlation between ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}e^{+}e^{-})$ (top panels) and between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}e^{+}e^{-})$ (bottom panels), for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The gray band corresponds to the upper bound ${\cal B}(\mu^{-}\to e^{-}e^{+}e^{-})<1\times 10^{-12}\,\,(\rm{at}\,90\%\,{\rm CL})$ [SINDRUM1988, Navas:PDG].
Figure 20: Correlation between ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-})$ due to conversion in nuclei (top panels) and between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-})$ due to conversion in nuclei (bottom panels), for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The gray band corresponds to the upper bound ${\cal B}(\mu^{-}\to e^{-})<4.3\times 10^{-12}\,\,(\rm{at}\,90\%\,{\rm CL})$[Kaulard:1998rb].
Figure 21: Correlation between ${\cal B}(\mu^{-}\to e^{-}e^{+}e^{-})$ and ${\cal B}(\mu^{-}\to e^{-}\gamma)$ (top panels) and between ${\cal B}(\mu^{-}\to e^{-}e^{+}e^{-})$ and ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})$ (bottom panels), for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The gray band corresponds to the upper bound ${\cal B}(\mu^{-}\to e^{-}e^{+}e^{-})<1\times 10^{-12}\,\,(\rm{at}\,90\%\,{\rm CL})$[SINDRUM1988, Navas:PDG].
Figure 22: Correlation between ${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ when constraints from LFV charged lepton decay modes are subsequently imposed as specified in the legenda, for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel).
4.2.3 Rare $B_{(s)}$ decays versus $\Delta a_{\mu}$
@@
As for the $Z^{\prime}$ contribution to the muon anomalous magnetic moment, $\Delta a_{\mu}$ vs ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ is shown in Fig. 23. $\Delta a_{\mu}$ turns out to be negative, it reaches at most $\Delta a_{\mu}\sim-10^{-12}$, a size similar to subleading hadron contributions such as the hadronic light-by-light contribution [Aoyama:2020ynm].
Figure 23: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and $\Delta a_{\mu}$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The light yellow points are selected requiring that the LFC ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ agrees with experiment at $1\sigma$.